# Upsample or Upweight? Balanced Training on Heavily Imbalanced Datasets

### v2

@@ 4.1 Setup
For our machine translation experiments, we use $\tau=5$ for the first 30k of training iterations and use $\tau=1$ for the second 30k. For our language modeling experiments, we use $\tau=5$ for the first 50k training iterations and $\tau=1$ for the second 50k. More detailed hyper-parameters are in Appendix D.
@@ 4.2 Main Results
EN IT ZH SW
Cooldown 2.59 2.63 2.56 2.94
Dense Update 1 3.17 2.89 2.85 3.11
Dense Update 2 3.31 3.02 3.35 3.16
Table 3: Validation loss on MC4 subset (lower is better) for different update schedules in Figure 5. Sparse updates perform better than dense updates.
@@
Table 4: SacreBLEU scores (higher is better) on a chosen subset of OPUS-100 with a mixture of high (1M), mid (500K - 1M), and low ($<$500K) resource languages. The best performance is bolded. Scores that are close (within 1 BLEU) of the best performance are colored in green. Scores lower than the best for more than 1.5 BLEU are highlighted in red. Our simple scheduling method outperforms various static and dynamic Temperature Sampling methods. We improve the performance on LRLs without sacrificing much performance on HRLs.
@@ 4.3 Findings / Ablations
We concluded that a decreasing schedule generally leads to faster convergence and better overall performance. Furthermore, we compare various fine-grained decreasing schedules that perform dense decreasing of the sampling temperature with our sparse update (5 for the first 50k, 1 for the second 50k training iterations) using the subset in mC4 described in Table 2. Figure 5 illustrates the schedules we compare, and 3 shows the validation loss of each decreasing schedule. Fine-grained online reweighting method yields worse performance, echoing the findings in Fan et al. (2024) and Xie et al. (2023). [^3] Both Xie et al. (2023) and Fan et al. (2024) approximate sampling probabilities using Scalarization when performing dense updates, which could also be the reason why dense update underperforms in their experiments.
@@ 5 Related Works
Gradient-based methods for Multi-Task Optimization
Training a multilingual model can be seen as Multi-Task Learning (MTL; Caruana (1997)), where each language is a single task. Gradient-based methods aim to reduce the discrepancy in directions of conflicting gradients in different tasks: PCGrad (Yu et al., 2020) aims to project the gradient of a task onto the orthogonal plane of the gradients of the other task, which explicitly aligns the gradient of different tasks. Gradient Vaccine (Wang et al., 2021) points out that using PCGrad to optimize multilingual pre-training relies on the false assumption that all languages are equally related and proposes to align the similarity of the gradient to the relatedness of the two languages. Another line of work (Wang et al., 2020a, b; Kreutzer et al., 2021) uses gradient similarity as the reward to train policy that decides the sampling probabilities for each domain in a Reinforcement Learning setting. Fan et al. (2024) utilizes gradient similarity between domains to design a temperature schedule for balancing multiple domains in training language models. However, recent studies point out that such gradient-based techniques do not yield significant improvement compared to a weighted sum of individual task losses (Scalarization) (Kurin et al., 2022; Xin et al., 2022; Royer et al., 2023). Our results also echo the findings of Zhai et al. (2023), where they find loss reweighting (Scalarization) underperforms standard training. Our work provides a possible explanation that Scalarization induces a larger variance in gradients.
@@
Another line of work utilizes the loss instead of the gradients per task for optimizing MTL models. An intuitive method is to put more weight on the task with the highest loss. In statistical learning, Distributionally Robust Optimization (DRO) methods (Ben-Tal et al., 2011; Duchi and Namkoong, 2021; Hashimoto et al., 2018; Sagawa* et al., 2020) instead of minimizing the sum of all losses, minimize the loss of the worst-performing subgroup to balance performance across all subgroups. Oren et al. (2019) and Xie et al. (2023) apply DRO to multi-domain language modeling to minimize the loss of a set of worse-performing domains. Similarly, Zhou et al. (2021) applies DRO to multilingual machine translation by minimizing the loss of a set of worse-performing translation pairs. One drawback of DRO-based methods is that they require training a proxy model on the dataset, which hinders their applicability on large-scale datasets. Cooldown can be seen as an efficient approximation of DRO methods by upsampling the worse-performing LRL and shifting the focus to the HRL once the LRL is sufficiently trained. Unlike DRO-based methods, Cooldown do does not require training proxy models (Liu et al., 2021) and dense updates on domain weights/sampling probabilities.
@@ Acknowledgements
This work is supported by a generous gift the Allen Institute for AI and partly by ONR grant (N00014-24-1-2089). We are grateful to Jingyu Zhang for his insightful feedback throughout this project. GPU machines for conducting experiments were provided by the DSAI cluster.
@@ Appendix B Proof of Theorem 2
Theorem 2 (Scalarization induces larger variance under Stochastic Gradient Descent) Using the same notation in Corollary 1.1, we have $\mathrm{Var}(\nabla\mathcal{L}_{\textrm{S}}(x))\geq\mathrm{Var}(\nabla\mathcal{L}_{\textrm{TS}}(x))$.
@@
Lemma 2.1 Let $\mathcal{D}=\{\mathcal{D}_{1},...,\mathcal{D}_{K}\}$, $\forall j\in\{1,...,K\}$, let $p(\mathcal{D}_{j},\tau)=\frac{\mathcal{D}_{j}^{\frac{1}{\tau}}}{\sum_{k}\mathcal{D}_{k}^{\frac{1}{\tau}}}$, then $\sum_{i=1}^{K}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}\geq 1$ holds true when $\forall j,\mathcal{D}_{j}\geq 0$.
@@
$\sum_{i=1}^{K}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}=\frac{\sum_{i}x_{i}^{\tau}}{(\sum_{i}x_{i})^{2}}\cdot\left(\sum_{i}x_{i}^{2-\tau}\right)$
@@
$1\leq\frac{\sum_{i}x_{i}^{\tau}}{(\sum_{i}x_{i})^{2}}\cdot\left(\sum_{i}x_{i}^{2-\tau}\right)=\sum_{i=1}^{K}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}.$
@@
Since $\nabla\mathcal{L}_{\textrm{S}}(x)$ and $\nabla\mathcal{L}_{\textrm{TS}}(x)$ are unbiased estimates of the total gradient, we only need to show that the expectation of the squared gradient is larger for the stochastic gradient under Scalarization.
$\underbrace{\mathop{\mathbb{E}}_{x\sim\mathcal{D}}[(w_{f(x)}\nabla\mathcal{L}(x))^{2}]}_{\text{Scalarization}}$ $=\sum_{i=1}^{K}p(\mathcal{D}_{i};1)w^{2}_{f(x)}\nabla\mathcal{L}^{2}(x)$
$\sum_{i=1}^{K}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}\nabla\mathcal{L}^{2}(x)$ $\geq\nabla\mathcal{L}^{2}(x)\sum_{i=1}^{K}p(\mathcal{D}_{i};\tau)\nabla\mathcal{L}^{2}(x)$
$=\underbrace{\mathop{\mathbb{E}}_{\mathcal{D}_{i}\sim p(\mathcal{D}_{i};\tau)}\left[\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{i}}[\nabla\mathcal{L}^{2}(x)]\right]}_{\text{Temperature Sampling}}.$
@@ Appendix C Proof of Theorem 3 and Construction of Figure 2
The difference in variance $\Delta=\mathrm{Var}(\nabla\mathcal{L}_{\textrm{S}}(x))-\mathrm{Var}(\nabla\mathcal{L}_{\textrm{TS}}(x))$ is non-decreasing when $\tau\geq 1$.
@@
$\sum_{i=1}^{K}\frac{p(\mathcal{D}_{i};\tau)^{2}}{p(\mathcal{D}_{i};1)}=\frac{\sum_{i}x_{i}^{\tau}}{(\sum_{i}x_{i})^{2}}\cdot\left(\sum_{i}x_{i}^{2-\tau}\right).$
@@
We plot the function $F(\tau)=\sum_{i}\frac{p(D_{i};\tau)^{2}}{p(D_{i};1)}$ against $\tau$ by starting with a uniform distribution and progressively increasing its skewness to resemble Zipf distributions. Specifically, we generate distributions $D_{i}\propto\frac{1}{i^{\alpha}}$ for various exponents $\alpha$ (ranging from 0 to higher values), where $i$ denotes the rank of each element. For each distribution, we compute the normalized probabilities $p(D_{i};\tau)=\frac{D_{i}^{1/\tau}}{\sum_{j}D_{j}^{1/\tau}}$ across a range of $\tau$ values. This approach allows us to analyze how increasing the skewness of the distribution influences the behavior of $F(\tau)$ as a function of $\tau$.

### v3

@@ 4.1 Setup
For our machine translation experiments, we use $\tau=5$ for the first 30k of training iterations and $\tau=1$ for the second 30k. For our language modeling experiments, we use $\tau=5$ for the first 50k training iterations and $\tau=1$ for the second 50k. Detailed hyper-parameters are in Appendix D.
@@ 4.2 Main Results
Table 3: SacreBLEU scores (higher is better) on a chosen subset of OPUS-100 with a mixture of high (1M), mid (500K - 1M), and low ($<$500K) resource languages. The best performance is bolded. Scores that are close (within 1 BLEU) of the best performance are colored in green. Scores lower than the best for more than 1.5 BLEU are highlighted in red. Cooldown outperforms various static and dynamic Temperature Sampling methods, by improving the performance on LRLs without sacrificing much performance on HRLs.
@@ 4.3 Findings / Ablations
Method EN IT ZH SW
Cooldown 2.59 2.63 2.56 2.94
Dense Update 1 3.17 2.89 2.85 3.11
Dense Update 2 3.31 3.02 3.35 3.16
Figure 6: Validation loss on mC4 subset (lower is better) for different update schedules in Figure 6. Sparse updates perform better than dense updates.
@@
We concluded that a decreasing schedule generally leads to faster convergence and better overall performance. Furthermore, we compare various fine-grained decreasing schedules that perform dense decreasing of the sampling temperature with our sparse update (5 for the first 50k, 1 for the second 50k training iterations) using the subset in mC4 described in Table 2. Figure 6 illustrates the schedules we compare, and 6 shows the validation loss of each decreasing schedule. Fine-grained online reweighting method yields worse performance, echoing the findings in Fan et al. (2024) and Xie et al. (2023). [^3] Both Xie et al. (2023) and Fan et al. (2024) approximate sampling probabilities using Scalarization when performing dense updates, which could also be the reason why dense update underperforms in their experiments.
@@ 5 Related Works
Gradient-based methods for Multi-Task Learning
Training a multi-domain language model can be seen as Multi-Task Learning (MTL; Caruana (1997)), where each domain is a single task. Gradient-based methods aim to reduce the discrepancy in directions of conflicting gradients in different tasks: PCGrad (Yu et al., 2020) aims to project the gradient of a task onto the orthogonal plane of the gradients of the other task. Another line of work (Wang et al., 2020a, b; Kreutzer et al., 2021) uses gradient similarity as the reward to train a policy that decides the sampling probabilities for each domain in a Reinforcement Learning setting. Fan et al. (2024) utilizes gradient similarity between domains to design a temperature schedule for balancing multiple domains in training language models. However, recent studies point out that such gradient-based techniques do not yield significant improvement compared to a weighted sum of individual task losses (Scalarization) (Kurin et al., 2022; Xin et al., 2022; Royer et al., 2023). Our results also echo the findings of Zhai et al. (2023), where they find loss reweighting (Scalarization) underperforms standard training. Our work provides a possible explanation that Scalarization induces a larger variance in gradients.
@@
Another line of work utilizes the loss instead of the gradients per task for optimizing MTL models. An intuitive method is to put more weight on the task with the highest loss. In statistical learning, Distributionally Robust Optimization (DRO) methods (Ben-Tal et al., 2011; Duchi and Namkoong, 2021; Hashimoto et al., 2018; Sagawa* et al., 2020) minimize the loss of the worst-performing subgroup to balance performance. Oren et al. (2019) and Xie et al. (2023) apply DRO to multi-domain language modeling to minimize the loss of a set of worse-performing domains. Similarly, Zhou et al. (2021) applies DRO to multilingual machine translation by minimizing the loss of a set of worse-performing translation directions. Cooldown can be seen as an efficient approximation of DRO methods by upsampling the worse-performing LRL and shifting the focus to the HRL once the LRL is sufficiently trained. Unlike DRO, Cooldown do does not require training proxy models (Liu et al., 2021) and dense updates on domain weights.
Scaling Laws for Domain Mixture
The search space for the optimal domain weights at any given training iteration is combinatorically large. Existing works conduct comprehensive experiments on smaller scaled models to learn how the training and generalization error varies according to dataset sizes and domain weights —— “scaling laws" of domain mixture Ye et al. (2024); ge2024bimixbivariatedatamixing; jiang2024adaptivedataoptimizationdynamic. Closer to our work, Chen et al. (2023) fits scaling laws for sampling temperature for multilingual machine translation. Concurrent to our work, he2024scalinglawsmultilinguallanguage fits scaling laws for multilingual language modeling. However, as jiang2024adaptivedataoptimizationdynamic denotes: “the optimal data policy for a smaller model does not necessarily generalize to larger models Albalak et al. (2023); Ye et al. (2024)."
@@ Acknowledgements
This work is supported by a generous gift the Allen Institute for AI and We are grateful Nicholas Lourie and Jingyu Zhang for their insightful feedback throughout this project. GPU machines for conducting experiments were provided by the DSAI cluster.
@@ Appendix B Proof of Theorem 2
Theorem 2 (Scalarization induces larger variance under Stochastic Gradient Descent) Using the same notation in Corollary 1.1, we have $\mathrm{Var}(\nabla\mathcal{L}_{S}(x;\mathbf{w}_{\tau}))\geq\mathrm{Var}(\nabla\mathcal{L}_{TS}(x;\tau))$.
@@
Lemma 2.1 Let $\mathcal{D}=\{\mathcal{D}_{1},...,\mathcal{D}_{K}\}$, $\forall j\in\{1,...,K\}$, let $p(j,\tau)=\frac{\mathcal{D}_{j}^{\frac{1}{\tau}}}{\sum_{k}\mathcal{D}_{k}^{\frac{1}{\tau}}}$, then $\sum_{i=1}^{K}\frac{p(i;\tau)^{2}}{p(i;1)}\geq 1$ holds true when $\forall j,\mathcal{D}_{j}\geq 0$.
@@
$\sum_{i=1}^{K}\frac{p(i;\tau)^{2}}{p(i;1)}=\frac{\sum_{i}x_{i}^{\tau}}{(\sum_{i}x_{i})^{2}}\cdot\left(\sum_{i}x_{i}^{2-\tau}\right)$
@@
$1\leq\frac{\sum_{i}x_{i}^{\tau}}{(\sum_{i}x_{i})^{2}}\cdot\left(\sum_{i}x_{i}^{2-\tau}\right)=\sum_{i=1}^{K}\frac{p(i;\tau)^{2}}{p(i;1)}.$
@@
Since $\nabla\mathcal{L}_{S}(x;\mathbf{w}_{\tau})$ and $\nabla\mathcal{L}_{TS}(x;\tau)$ are unbiased estimates of the total gradient, we only need to show that the expectation of the squared gradient is larger for the stochastic gradient under Scalarization.
$\underbrace{\mathop{\mathbb{E}}_{x\sim\mathcal{D}}[(w_{f(x)}\nabla\mathcal{L}(x))^{2}]}_{\text{Scalarization}}$ $=\sum_{i=1}^{K}p(i;1)w^{2}_{f(x)}\nabla\mathcal{L}^{2}(x)$
$\sum_{i=1}^{K}\frac{p(i;\tau)^{2}}{p(i;1)}\nabla\mathcal{L}^{2}(x)$ $\geq\nabla\mathcal{L}^{2}(x)\sum_{i=1}^{K}p(i;\tau)\nabla\mathcal{L}^{2}(x)$
$=\underbrace{\mathop{\mathbb{E}}_{\mathcal{D}_{i}\sim p(i;\tau)}\left[\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{i}}[\nabla\mathcal{L}^{2}(x)]\right]}_{\text{Temperature Sampling}}.$
@@ Appendix C Proof of Theorem 3 and Construction of Figure 2
The difference in variance $\Delta=\mathrm{Var}(\nabla\mathcal{L}_{S}(x))-\mathrm{Var}(\nabla\mathcal{L}_{TS}(x))$ is non-decreasing when $\tau\geq 1$.
@@
$\sum_{i=1}^{K}\frac{p(i;\tau)^{2}}{p(i;1)}=\frac{\sum_{i}x_{i}^{\tau}}{(\sum_{i}x_{i})^{2}}\cdot\left(\sum_{i}x_{i}^{2-\tau}\right).$
@@
We plot the function $F(\tau)=\sum_{i}\frac{p(i;\tau)^{2}}{p(D_{i};1)}$ against $\tau$ by starting with a uniform distribution and progressively increasing its skewness to resemble Zipf distributions. Specifically, we generate distributions $D_{i}\propto\frac{1}{i^{\alpha}}$ for various exponents $\alpha$ (ranging from 0 to higher values), where $i$ denotes the rank of each element. For each distribution, we compute the normalized probabilities $p(i;\tau)=\frac{\mathcal{|}{D}_{i}|^{1/\tau}}{\sum_{j}|\mathcal{D}_{j}|^{1/\tau}}$ across a range of $\tau$ values. This approach allows us to analyze how increasing the skewness of the distribution influences the behavior of $F(\tau)$ as a function of $\tau$.